# FastCuRL: Curriculum Reinforcement Learning with Stage-wise Context Scaling for Efficient Training R1-like Reasoning Models

### v4

@@ Abstract
Improving training efficiency continues to be one of the primary challenges in large-scale Reinforcement Learning (RL). In this paper, we investigate how context length and the complexity of training data influence the RL scaling training process of R1-distilled small reasoning models, e.g., DeepSeek-R1-Distill-Qwen-1.5B. Our experimental results reveal that: (1) simply controlling the context length and curating the training data based on the input prompt length can effectively improve the training efficiency of scaling RL, achieving better performance with more concise CoT; (2) properly scaling the context length helps mitigate entropy collapse; and (3) choosing an optimal context length can improve the efficiency of model training and incentivize the model’s chain-of-thought reasoning capabilities. Inspired by these insights, we propose FastCuRL, a curriculum RL framework with stage-wise context scaling to achieve efficient training and concise CoT reasoning. Experiment results demonstrate that FastCuRL-1.5B-V3 significantly outperforms state-of-the-art reasoning models on five competition-level benchmarks and achieves 49.6% accuracy on AIME 2024. Furthermore, FastCuRL-1.5B-Preview surpasses DeepScaleR-1.5B-Preview on five benchmarks while only using a single node with 8 GPUs and a total of 50% of training steps. The code, training data, and models have been publicly released111https://github.com/nick7nlp/FastCuRL.
@@ 1 Introduction
A primary finding from recent breakthroughs, exemplified by DeepSeek-R1 DeepSeek-AI (2025), reveals a scaling phenomenon in the training process of Reinforcement Learning (RL). Inspired by these findings, training LLMs through scaling RL has recently emerged as a promising paradigm for addressing complex reasoning tasks and many valuable research endeavours Luo et al. (2025); Face (2025); Hu et al. (2025); Zeng et al. (2025); Liu et al. (2025) have emerged to explore and replicate reasoning models akin to DeepSeek-R1 (for example, starting from R1-distilled or pre-trained models) by extending the generation length of CoT.
However, excessively long CoT responses significantly increase computational overhead during model training and deployment. Moreover, recent studies Yeo et al. (2025); Wu et al. (2025); Team (2025a) have identified an inherent overthinking phenomenon in reasoning models, which includes irrelevant details and repetitive thinking patterns. This kind of information leads to inefficient use of computational resources and undermines reasoning accuracy, which causes models to stray from valid logical pathways, resulting in incorrect answers.
To this end, recent studies Team (2025a); Luo et al. (2025); Liu et al. (2025); Yu et al. (2025) focus on efficient reasoning for optimizing the model to generate more concise solutions. Among them, DeepScaleR Luo et al. (2025) propose to iteratively increase the context length from 8K to 24K to train the DeepSeek-R1-Distill-Qwen-1.5B model toward more concise reasoning, outperforming OpenAI’s o1-preview OpenAI (2024). By observing the training logs222https://github.com/agentica-project/rllm of DeepScaleR in Figure 2, we find two issues:
When the context length is 8K, about 42% of the model’s outputs are clipped, which reduces the model’s training efficiency.
When the context length is 24K, the model’s entropy collapses. Entropy reflects the exploration capability of an LLM during training. A rapid decrease in entropy might lead to premature convergence, preventing the model from achieving the expected performance.
@@ 2 Methodology
In this section, we introduce our investigation into how the model’s context length and the complexity of training data influence the training process of R1-like reasoning models. Specifically, our method consists of two main components: (1) curating a complexity-aware, mathematics-focused dataset, and (2) implementing a resource-efficient reinforcement learning algorithm. These two components aim to balance a trade-off between achieving performance improvements and addressing practical limitations, such as reducing computational costs.
Example Problem (Output Length=74706 characters): Ashley, Betty, Carlos, Dick, and Elgin went shopping. Each had a whole number of dollars to spend, and together they had dollars. The absolute difference between the amounts Ashley and Betty had to spend was dollars. The absolute difference between the amounts Betty and Carlos had was dollars, between Carlos and Dick was dollars, between Dick and Elgin was dollars, and between Elgin and Ashley was dollars. How many dollars did Elgin have? |
2.1 Complexity-Aware Data Curation
To ensure a fair comparison, we directly employ the dataset from DeepScaleR as the training data. The DeepScaleR dataset Luo et al. (2025) consists of 40,315 unique mathematics-specific problem-answer pairs collected from AIME (1984-2023), AMC (prior to 2023), Omni-MATH, and the Still dataset Balunović et al. (2025); Gao et al. (2024); Min et al. (2024). The statistics of the DeepScaleR dataset are shown in Figure 3.
As illustrated in Figure 2, over 42% of training samples are clipped at the beginning of the training steps due to exceeding the maximum response length. By observing and analyzing the clipped responses, we find that they mainly correspond to two types of problems. The first type pertains to challenging problems requiring long CoT responses to solve. The second involves questions laden with numerous conditions, prompting the model to verify each condition repeatedly during problem-solving, e.g., the problem shown in Table 1. This repetitive verification may result in redundant thinking patterns, ultimately causing the reasoning responses to be unduly long. Both situations may impact the model’s training efficiency during the 8K context.
Datasets | Average Input Prompt Length |
L1 | 148.65 |
L2 | 247.24 |
L3 | 407.78 |
@@ 3.1 Experimental Setup
Experiments | Stages | Context Length | Training Data | Batch Size | Rollout | Avg. | |||
Input | Output (K) | (for Entropy) | (for KL) | ||||||
Exp-1 | 3 | 1K | 8, 16, 24 | 1, 2, 3 | 128, 64, 64 | 8, 8, 8 | 0.001 | 0.001 | 0.550 |
Exp-2 | 3 | 1K | 8, 16, 24 | 1, 3, 2 | 128, 64, 64 | 0.540 | |||
Exp-3 | 3 | 1K | 8, 16, 24 | 1, 2, 2 | 128, 64, 64 | 0.552 | |||
Exp-4 | 4 | 1K | 8, 16, 24, 32 | 1, 2, 3, 2 | 128, 64, 64, 64 | 8, 8, 8, 16 | 0.001 | 0.001 | 0.566 |
Exp-5 | 4 | 1K | 8, 16, 24, 24 | 1, 2, 3, 2 | 128, 64, 64, 64 | 0.565 | |||
Exp-6 | 4 | 1K | 8, 16, 24, 16 | 1, 2, 3, 2 | 128, 64, 64, 64 | 0.575 | |||
Exp-7 | 5 | 1K | 8, 16, 24, 16, 24 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.001 | 0.001 | 0.556 |
Exp-8 | 5 | 1K | 8, 16, 24, 16, 16 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 0.567 | |||
Exp-9 | 5 | 1K | 8, 16, 24, 16, 8 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 0.535 | |||
Exp-10 | 5 | 1K | 8, 16, 24, 16, 16 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.000001 | 0.000 | 0.600 |
Exp-11 | 5 | 1K | 8, 16, 24, 16, 16 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.000 | 0.000 | 0.616 |
@@ 3.3 Benchmarks
To comprehensively evaluate the performance, we select five competition-level benchmark datasets: MATH 500 Hendrycks et al. (2021), AIME 2024333https://huggingface.co/datasets/AI-MO/aimo-validation-aime, AMC 2023444https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Minerva Math Lewkowycz et al. (2022), and OlympiadBench He et al. (2024).
@@ 3.5.1 Dataset Complexity Verification
Model | MATH 500 | AIME 2024 | AMC 2023 | Minerva Math | OlympiadBench | Avg. |
@@ 3.5.2 Multi-Stage Experimental Results
In the third set of experiments, we find that neither increasing nor decreasing the context length is as effective as maintaining the context length at 16K. Does this phenomenon suggest that there is a "sweet spot" for context length in R1-like models, and that for the DeepSeek-R1-Distill-Qwen-1.5B, 16K is the optimal sweet spot? Or is it that 16K is closer to the sweet spot compared to 24K and 8K? Based on this question, we conduct a series of experiments where we train the model with different context lengths and set the entropy coefficient equal to to observe the changes in the entropy. As shown in Figure 7, we find that when the context lengths are 4K, 8K, and 12K, the entropy rapidly decreases to a small value, indicating that the model has lost its exploratory capability. Interestingly, when the context lengths are 16K, 20K, and 24K, the entropy stabilizes at a fixed value and does not decrease rapidly.
@@ 3.5.3 Overall Comparison Results
Total | Correct | Incorrect | Total | Correct | Incorrect | |
DeepSeek-R1-Distill-Qwen-1.5B | 43176 | 21859 | 52629 | 109 | 49 | 138 |
FastCuRL-1.5B-Preview | 28681 | 18970 | 36044 | 80 | 48 | 104 |
3.5.4 The Effectiveness of KL and Entropy
The KL penalty and entropy loss are very important in RL training. Therefore, we conduct simple ablation experiments on the KL penalty and entropy loss. As presented in Figure 8, we find that when training the DeepSeek-R1-Distill-Qwen-1.5B model without the KL penalty, even when the average output length was compressed to between 3500-4000 tokens, the model’s output length does not show a significant increasing trend. From the results in Figure 9, we can see that removing the entropy loss caused the model’s output length to decrease significantly around step 800. In Figure 8 and Figure 9, the blue lines represent the original experimental setup, but these are results from two different experiments. This paper primarily focuses on exploring the impact of context length and data complexity on the training process. Therefore, we do not provide an extensive analysis of the effects of the KL penalty and entropy loss.

### v5

@@ Abstract
Improving training efficiency continues to be one of the primary challenges in large-scale Reinforcement Learning (RL). In this paper, we investigate how context length and the complexity of training data influence the RL scaling training process of R1-distilled reasoning models, e.g., DeepSeek-R1-Distill-Qwen-1.5B. Our experimental results reveal that: (1) simply controlling the context length and curating the training data based on the input prompt length can effectively improve the training efficiency of RL scaling, achieving better performance with more concise CoT; (2) properly scaling the context length helps mitigate entropy collapse; and (3) carefully choosing the context length facilitates achieving efficient LLM training and reasoning. Inspired by these insights, we propose FastCuRL, a curriculum RL framework with stage-wise context scaling to achieve efficient LLM training and reasoning. Extensive experimental results demonstrate that FastCuRL-1.5B-V3 significantly outperforms state-of-the-art reasoning models on five competition-level benchmarks and achieves 49.6% accuracy on AIME 2024. Furthermore, FastCuRL-1.5B-Preview surpasses DeepScaleR-1.5B-Preview on five benchmarks while only using a single node with 8 GPUs and a total of 50% of training steps. The code, training data, and models have been publicly released111https://github.com/nick7nlp/FastCuRL.
@@ 1 Introduction
A primary finding from recent breakthroughs, exemplified by DeepSeek-R1 DeepSeek-AI (2025), reveals a scaling phenomenon in the training process of Reinforcement Learning (RL). Inspired by these findings, training LLMs through scaling RL has recently emerged as a promising paradigm for addressing complex reasoning tasks and many valuable research endeavours Face (2025); Hu et al. (2025); Zeng et al. (2025); Liu et al. (2025) have emerged to explore and replicate reasoning models akin to DeepSeek-R1 (for example, starting from R1-distilled or pre-trained models) by extending the generation length of CoT.
However, excessively long CoT responses significantly increase computational overhead during model training and deployment. Moreover, recent studies Yeo et al. (2025); Wu et al. (2025); Team (2025a) have identified an inherent overthinking phenomenon in reasoning models, which includes irrelevant details and repetitive thinking patterns. This kind of information leads to inefficient use of computational resources and undermines reasoning accuracy, which causes models to stray from valid logical pathways, resulting in incorrect answers. To this end, recent studies Team (2025a); Luo et al. (2025); Liu et al. (2025); Yu et al. (2025) focus on efficient reasoning for optimizing the model to generate more concise solutions. Among them, DeepScaleR Luo et al. (2025) propose to iteratively increase the context length from 8K to 24K to train the DeepSeek-R1-Distill-Qwen-1.5B model toward more concise reasoning, outperforming OpenAI’s o1-preview OpenAI (2024). By observing the training logs222https://github.com/agentica-project/rllm of DeepScaleR in Figure 2, we find two issues:
When the context length is 8K, about 42% of the model’s outputs are clipped, which reduces the training efficiency.
When training in the third stage with a context length of 24K, the model’s entropy collapses333Entropy reflects the exploration capability of LLMs during RL training. Typically, a rapid decrease in entropy might lead to premature convergence, preventing the model from achieving the expected performance., reducing its exploration capabilities.
@@ 2 Methodology
In this section, we introduce our method in detail, which consists of three main parts: (1) a resource-efficient RL algorithm, (2) a complexity-aware, mathematics-focused dataset, and (3) a stage-wise context scaling approach. These three components aim to achieve efficient LLM reasoning.
@@ 2.1 Reinforcement Learning Algorithm
2.2 Complexity-Aware Data Curation
To ensure a fair comparison, we directly employ the dataset from DeepScaleR as the training data. The DeepScaleR dataset Luo et al. (2025) consists of 40,315 unique mathematics-specific problem-answer pairs collected from AIME (1984-2023), AMC (prior to 2023), Omni-MATH, and the Still dataset Balunović et al. (2025); Gao et al. (2024); Min et al. (2024). We perform a simple filtering of the original dataset based on prompt length, resulting in 40,207 data samples. The statistics of the filtered dataset are shown in Figure 3.
Data | Prompt Len. | Output Len. | #Samples |
L1 | 48.26 | 8805.78 | 25168 |
L2 | 78.03 | 10063.44 | 40207 |
L3 | 132.22 | 12168.16 | 15039 |
As illustrated in Figure 2, during the first training stage, over 42% of training samples are clipped at the beginning of the training steps due to exceeding the maximum response length. By observing and analyzing the clipped responses, we find that they mainly correspond to two types of problems. The first type pertains to challenging problems requiring long CoT responses to solve. The second involves questions laden with numerous conditions, prompting the model to verify each condition repeatedly during problem-solving. This repetitive verification may result in redundant thinking patterns, ultimately causing the reasoning responses to be unduly long. Both situations may impact the model training efficiency during the 8K context.
@@ 3.1 Experimental Setup
Experiments | Training | Context Length | Training Data | Batch Size | Rollout | Avg. | |||
(Stages) | Steps | Input | Output (K) | (for Entropy) | (for KL) | ||||
Exp-1 (3) | 3 | 1K | 8, 16, 24 | 1, 2, 3 | 128, 64, 64 | 8, 8, 8 | 0.001 | 0.001 | 0.550 |
Exp-2 (3) | 3 | 1K | 8, 16, 24 | 1, 3, 2 | 128, 64, 64 | 8, 8, 8 | 0.001 | 0.001 | 0.540 |
Exp-3 (3) | 3 | 1K | 8, 16, 24 | 1, 2, 2 | 128, 64, 64 | 8, 8, 8 | 0.001 | 0.001 | 0.552 |
Exp-4 (4) | 4 | 1K | 8, 16, 24, 32 | 1, 2, 3, 2 | 128, 64, 64, 64 | 8, 8, 8, 16 | 0.001 | 0.001 | 0.566 |
Exp-5 (4) | 4 | 1K | 8, 16, 24, 24 | 1, 2, 3, 2 | 128, 64, 64, 64 | 8, 8, 8, 16 | 0.001 | 0.001 | 0.565 |
Exp-6 (4) | 4 | 1K | 8, 16, 24, 16 | 1, 2, 3, 2 | 128, 64, 64, 64 | 8, 8, 8, 16 | 0.001 | 0.001 | 0.575 |
Exp-7 (5) | 5 | 1K | 8, 16, 24, 16, 8 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.001 | 0.001 | 0.535 |
Exp-8 (5) | 5 | 1K | 8, 16, 24, 16, 16 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.001 | 0.001 | 0.567 |
Exp-9 (5) | 5 | 1K | 8, 16, 24, 16, 24 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.001 | 0.001 | 0.556 |
Exp-10 (5) | 5 | 1K | 8, 16, 24, 16, 16 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.000001 | 0.000 | 0.600 |
Exp-11 (5) | 5 | 1K | 8, 16, 24, 16, 16 | 1, 2, 3, 2, 2 | 128, 64, 64, 64, 64 | 8, 8, 8, 16, 16 | 0.000 | 0.000 | 0.616 |
@@ 3.3 Benchmarks
To comprehensively evaluate the performance, we select five competition-level benchmark datasets: MATH 500 Hendrycks et al. (2021), AIME 2024444https://huggingface.co/datasets/AI-MO/aimo-validation-aime, AMC 2023555https://huggingface.co/datasets/AI-MO/aimo-validation-amc, Minerva Math Lewkowycz et al. (2022), and OlympiadBench He et al. (2024).
@@ 3.5.1 Dataset Complexity Verification
Model | MATH 500 | AIME 2024 | AMC 2023 | Minerva Math | OlympiadBench | Avg. Score |
@@ 3.5.2 Multi-Stage Experimental Results
In the third set of experiments, we observe that neither increasing nor decreasing the context length from 16K yields comparable effectiveness, suggesting that maintaining the context length at 16K produces optimal results. This phenomenon raises an important question: Does there exist an optimal "sweet spot" for context length in R1-like models? Specifically, for the DeepSeek-R1-Distill-Qwen-1.5B model, is 16K the precise optimal context length, or does it merely approximate the optimal value more closely than the alternative configurations of 8K and 24K?
To investigate this question, we conducted a comprehensive series of experiments examining the model’s behavior across various context lengths. We trained the model with different context length configurations while maintaining the entropy coefficient at to observe entropy dynamics. As illustrated in Figure 6, our findings reveal a notable dichotomy in model behavior. For context lengths of 4K, 8K, and 12K, the entropy exhibits a rapid decline, suggesting a significant degradation in the model’s exploratory capabilities. In contrast, when the context lengths are configured at 16K, 20K, and 24K, the entropy demonstrates remarkable stability, converging to and maintaining a consistent value.
Total | Correct | Incorrect | Total | Correct | Incorrect | |
DeepSeek-R1-Distill-Qwen-1.5B | 16390 | 8126 | 18775 | 107 | 49 | 129 |
FastCuRL-1.5B-Preview | 8654 | 7091 | 12010 | 80 | 47 | 104 |